Derive the expression for moment of inertia of a rod about it's centre and perpendicular to the rod

Grade:11

1 Answers

k sibin
5 years ago
lets take a dm mass x distance from its centre on the rod of mass m and length l
I=∫l/2-l/2 (m/l).dx)x2
 
 =m/l[x3/3]-l/2l/2 
=m/l[l3/24 + l3/24]
=ml/12